OpenStudy (anonymous):

A wedding cake is prepared using 3 individual cakes. The first is a rectangular prism with a length thirteen inches, a width 9 inches, and a height 3 inches. The second is a cylinder with a diameter 6 inches and a height 4 inches. The third is a cube with side length 3 inches. What is the approximate total volume of the cake? Use three point one four for pi. Round to the nearest whole number.

OpenStudy (ash2326):

We have three cakes, Let's calculate the volume of each individually. The volume of cake will be the sum of the volumes of the individual cakes. Let's get to work:)

OpenStudy (ash2326):

The first is a rectangular prism with length 13 inches, width =9 inches and height = 3 inches l= 13 inches w= 9 inches h= 3 inches  Volume of this rectangular prism= \(l\times w\times h\) =\(13\times 9 \times 3= 351\ inches^3\)